# Geodesic Slice Sampling on the Sphere

### v3

@@ 3.1 Ideal Geodesic Slice Sampling
1:  Draw $V\sim\mathcal{U}_{\mathbb{S}^{d-2}_{x}}$, call the result $v$.   {Perform Algorithm 5 with input $x$.}
@@
The geodesic slice sampler can be viewed as an instance of the procedure described in Section 2.3 where the second step uses a distribution on the geodesic that depends on the initial point $x$ and the chosen geodesic $\gamma$.
@@
The right-hand side of (7) depends on $p$ by
@@
Observe that the decreasing level set function $t\mapsto\mathcal{U}_{\mathbb{S}^{d-1}}(L(t))$ determines crucially the size of $r_{p}\in(0,1]$, cf. Roberts and Rosenthal (1999); Natarovskii et al. (2021b) where it also appears differently in the characterization of convergence statements of slice sampling.
@@
Proof (Proof of Theorem 6.) Let $x\in\mathbb{S}^{d-1}$ and $s\in(0,\left\|p\right\|_{\infty})$ such that $L(s)$ is open and non-empty. Observe that for all $v\in{\mathbb{S}^{d-2}_{x}}$ and $\theta\in[0,2\uppi)$ with $\gamma_{(x,v)}(\theta)\in L(s)$ holds
@@
for any $s>0$. Taking the infimum over $s>0$ on the right hand-side gives (7). By the lower semicontinuity of $p$ we have that $L(s)$ is open and non-empty, such that $\mathcal{U}_{\mathbb{S}^{d-1}}(L(s))\in(0,1]$ for all $s\in(0,\left\|p\right\|_{\infty})$. This implies $0<s\;\mathcal{U}_{\mathbb{S}^{d-1}}(L(s))<\left\|p\right\|_{\infty}$, which provides the final claim.
@@
Now we modify the ideal geodesic slice sampler by replacing the acceptance/rejection step of Algorithm 2 by a shrinkage procedure that has been also used in elliptical slice sampling (see Murray et al., 2010) and originates in Neal’s bracketing procedure on the interval $[0,2\uppi)$ (see Neal, 2003). We pick candidates from subsets of $[0,2\uppi)$ that shrink until an accepted sample is generated. Intuitively, this strategy reduces the number of rejections per iteration, because candidates will be drawn from nested intervals that contain a neighborhood of the current state. Note that this approach can still be viewed as a realization of the procedure discussed in Section 2.3 in the sense of Remark 4.
@@
A single transition of the geodesic shrinkage slice sampler is presented in Algorithm 4 and for an illustration of the shrinkage scheme on a geodesic on $\mathbb{S}^{2}$ we refer to Figure 2.
@@
1:  Draw $V\sim\mathcal{U}_{\mathbb{S}^{d-2}_{x}}$, call the result $v$.   {Perform Algorithm 5 with input $x$.}
@@
3:  Draw $\Theta\sim Q_{{x,}{v,}{t}}$, call the result $\theta$.   {Perform Algorithm 3 with input $x,v,t$.}
@@
Observe that due to the $2\uppi$-periodicity of the geodesics on the sphere, we have
@@
Proof (Proof of Theorem 16.) For arbitrary $x\in\mathbb{S}^{d-1}$, $s\in(0,\left\|p\right\|_{\infty})$ with $L(s)$ being open and non-empty, for any $v\in{\mathbb{S}^{d-2}_{x}}$ and $\theta\in[0,2\uppi)$ with $\gamma_{(x,v)}(\theta)\in L(s)$ we have (cf. (8)) that
@@
Moreover, for all $t\in(0,p(x))$ we obtain with $\Theta\sim\mathcal{U}_{(0,2\uppi)}$ that
$Q_{{x,}{v,}{t}}(B)\geq\mathbb{P}(\Theta\in B\cap L(x,v,t))=\frac{1}{2\uppi}{\lambda}(B\cap L(x,v,t)),\quad B\in\mathcal{B}([0,2\uppi)).$ (11)
@@
Note that the right hand-side of (9) coincides with the right hand-side of (7) stated in Theorem 6 that addressed the uniform ergodicity statement for ideal slice sampling. This is due to the proof technique, since in the small set estimate, we lower bound both kernels with the same expression of Lemma 11. Intuitively, it is clear that $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(H^{n}(x,\cdot),\pi)$ is smaller than $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(\widetilde{H}^{n}(x,\cdot),\pi)$, since the shrinkage procedure just adaptively imitates the acceptance/rejection step of the ideal one to gain computational efficiency. Exactly this gain in efficiency leads in applications to a potentially better accuracy to cost ratio, although the performance per Markov chain transition may be worse. We also point out that essentially Remark 7 and Remark 8 apply for Theorem 16 as well.
@@ 4 Numerical Illustrations
We demonstrate geodesic slice sampling on the sphere for approximate sampling on various distributions on the sphere and compare it to tailor-made and general purpose MCMC algorithms including random-walk Metropolis-Hastings (RWMH) and Hamiltonian Monte Carlo (HMC). The RWMH algorithm on the sphere is very similar to the standard RWMH, for details refer to Appendix F. Lan et al. (2014) have developed a modified version of HMC for sampling spherical distributions. The details of the HMC algorithm can be found in Appendix G. Both RWMH and HMC have a step-size parameter that is automatically tuned for each target, see Appendix H. For a prespecified sampler we denote by $(x_{n})_{n\in\mathbb{N}}\subset\mathbb{S}^{d-1}$ a realization of the corresponding Markov chain.
4.1 Bingham Distribution
In our first test, we aim to sample from the Bingham distribution (Mardia and Jupp, 2000) whose unnormalized density is defined as
$p_{\text{Bing}}{}(x)=\exp(x^{T}\!Ax)$ (12)
for $x\in\mathbb{S}^{d-1}$. Without loss of generality, the matrix $A\in\mathbb{R}^{d\times d}$ is symmetric such that we can transform the variables into the eigenbasis of $A$. In the following, we assume that the variables underwent this transformation by letting $A=\text{diag}(\kappa_{1},\ldots,\kappa_{d})$ where $\kappa_{1}\leq\kappa_{2}\leq\ldots\leq\kappa_{d}$ are the eigenvalues of $A$. Because the Bingham distribution is invariant under shifts of the diagonal of $A$, i.e., invariant under $A\mapsto A+c\,\mathrm{Id}_{d}$, where $c\in\mathbb{R}$ is some constant, we can let $\kappa_{1}=0$. The Bingham distribution is bimodal with symmetric modes at ±$\pm u_{d}$ where $u_{d}\in\mathbb{S}^{d-1}$ is the eigenvector of $A$ with the largest eigenvalue $\kappa_{d}$. The maximal value of the logarithm of the unnormalized probability density is $\kappa_{d}$, i.e., $\max_{x\in\mathbb{S}^{d-1}}\log p_{\text{Bing}}{}(x)=\kappa_{d}$. Various algorithms for simulating the Bingham distribution have been proposed. The algorithm by Kent et al. (2018) is particularly attractive because it uses an acceptance/rejection sampler based on the angular central Gaussian (ACG) distribution as an envelope and is therefore straightforward to implement. We use this method as a baseline against which we compare our slice samplers as well as RWMH and HMC. We run the geodesic slice samplers on the sphere (geoSSS) as well as RWMH and HMC on two high-dimensional Bingham targets with $d=10$ and $d=50$, respectively. For each target and sampling algorithm, we simulate the consecutive Markov chain realizations $x_{1},\dots,x_{N}$ with $N=10^{5}$ starting from the mode $u_{d}$. To explore the variability of all MCMC algorithms, we run 10 repetitions. The step-size parameter of the RWMH and HMC algorithm is tuned as described in Appendix H to achieve a reasonable average acceptance rate.
Figure 3: Traces of approximate samples from a Bingham distribution with $d=10$ and $\kappa_{d}=30$ projected onto one of both modes $u_{d}$.
The first Bingham target has $d=10$ and $\kappa_{d}=30$. In Figure 3, we show traces of approximate Bingham samples projected onto the first mode, that is $u_{d}^{T}x$, which varies between $1$ and $-1$ and is expected to peak somewhat below these extreme values that correspond to both modes. As is evident from the trace plots, both variants of geodesic slice sampling on the sphere find both modes of the Bingham distribution and rapidly jump between them. RWMH, on the other hand, does not escape from the mode in which the chain was started and therefore only finds a single mode. HMC also finds the second mode, but only jumps occasionally between the two modes.
Figure 4: Histograms of approximate Bingham samples with $d=10$ and $\kappa_{d}=30$ projected on the first mode obtained with each MCMC method are shown in grey. The red dashed line indicates the baseline obtained with the acceptance/rejection sampler of Kent et al. (2018).
The insufficient sampling by RWMH and HMC results in an incorrect exploration of the modes, which should be populated equally. This is illustrated in Figure 4 which also shows the distribution of $u_{d}^{T}x$ obtained by the aforementioned acceptance/rejection method (see Kent et al., 2018). The histograms obtained with the geodesic slice samplers closely match the baseline, whereas RWMH completely misses the second mode and HMC misrepresents the probability mass under the modes.
Figure 5: Left: Autocorrelation analysis of $x\mapsto u_{d}^{T}x$ with respect to approximate samples from a Bingham distribution with $d=10$ and $\kappa_{d}=30$. Right: Estimated hopping frequency between both modes of the Bingham distribution. Note that the hopping frequency is shown on a logarithmic axis.
Figure 5 shows the autocorrelation function (ACF) of $u_{d}^{T}x$ for the four MCMC samplers. We see a rapid decorrelation in case of both geodesic slice samplers; in fact, the ACF of rejection-based geoSSS drops to zero after a single step. Samples generated with RWMH decorrelate after roughly $200$ MCMC steps, but the faster decorrelation (in comparison to HMC) is due to the fact that RWMH only explores a single mode. HMC finds both modes but shows a very slowly decaying ACF, because jumps between the two modes only occur very rarely. This is also reflected in the effective sample size (ESS). Due to the immediate decorrelation of samples generated with geoSSS using a rejection strategy, the relative ESS [^1] The relative ESS is just the ESS divided by the total number of performed MCMC steps, in our experiments $10^{5}$. is estimated to be 99.73 %. The shrinkage-based geoSSS obtains a relative ESS of 15.2 %, whereas RMWH and HMC achieve only very low relative ESS: 0.004 % and 0.01 %, respectively.
To quantify how rapidly the MCMC samplers mix between the two modes of the Bingham distribution, we estimated a hopping frequency, which we define as the average number of times the Markov chain jumps from one mode to the other, given by
$\frac{1}{N-1}\sum_{n=1}^{N-1}\llbracket\text{sign}(u_{d}^{T}x_{n+1})\not=\text{sign}(u_{d}^{T}x_{n})\rrbracket,$ (13)
[…]
Figure 5 shows the estimated hopping frequencies. As expected, the rejection-based geoSSS shows the highest number of oscillations between both modes with approximately 50% hopping frequency, whereas the shrinkage-based geoSSS tends to jump only every seventh step to the other mode. This behavior is expected, because the geodesic level set always contains both modes with equal probability. Therefore, the rejection sampling strategy finds each mode with equal probability, independent of what the current state of the Markov chain is. The shrinkage-based approach has a higher chance to stay in the vicinity of the current state.
Figure 6: Geodesic distance (log-scale) between successive approximate samples for a Bingham distribution with $d=10$ and $\kappa_{d}=30$.

### v4

@@ 3.1 Ideal Geodesic Slice Sampling
1: Draw $V\sim\mathcal{U}_{\mathbb{S}^{d-2}_{x}}$, call the result $v$.   {Perform Algorithm 5 with input xx.}
@@
The geodesic slice sampler can be viewed as an instance of the procedure described in Section 2.3 where the second step uses a distribution on the great circle that depends on the initial point $x$ and the chosen great circle $\gamma$.
@@
The right-hand side of (7) depends on $p$ through
@@
Observe that the size of $r_{p}\in(0,1]$ solely depends on the decreasing level set function $t\mapsto\mathcal{U}_{\mathbb{S}^{d-1}}(L(t))$. This function plays an even stronger role in controlling the convergence of ideal slice sampling. It is known to be the only relevant quantity for geometric ergodicity and spectral gap estimates, cf. Roberts and Rosenthal (1999); Natarovskii et al. (2021b).
@@
Proof (Proof of Theorem 6.) Let $x\in\mathbb{S}^{d-1}$ and $s\in(0,\left\|p\right\|_{\infty})$. Observe that for all $v\in{\mathbb{S}^{d-2}_{x}}$ and $\theta\in[0,2\uppi)$ with $\gamma_{(x,v)}(\theta)\in L(s)$ holds
@@
for any $s\in(0,\|p\|_{\infty})$. For $s\geq\|p\|_{\infty}$ the statement holds trivially. Therefore taking the infimum over $s>0$ on the right hand-side gives (7). By the lower semicontinuity of $p$ we have that $L(s)$ is open and non-empty, such that $\mathcal{U}_{\mathbb{S}^{d-1}}(L(s))\in(0,1]$ and consequently $0<s\;\mathcal{U}_{\mathbb{S}^{d-1}}(L(s))<\left\|p\right\|_{\infty}$ for all $s\in(0,\left\|p\right\|_{\infty})$, which provides the final claim.
@@
Now we modify the ideal geodesic slice sampler by replacing the acceptance/rejection step of Algorithm 2 by a shrinkage procedure that has been also used in elliptical slice sampling (see Murray et al., 2010) and originates in Neal’s bracketing procedure on the interval $0,2\uppi)$ (see Neal, 2003). We pick candidates from subsets of $0,2\uppi)$ that shrink until an accepted sample is generated. Intuitively, this strategy reduces the number of rejections per iteration, because candidates will be drawn from nested intervals that contain a neighborhood of the current state. Note that this approach can still be viewed as a realization of the procedure discussed in Section 2.3 in the sense of Remark 4.
@@
A single transition of the geodesic shrinkage slice sampler is presented in Algorithm 4 and for an illustration of the shrinkage scheme on a great circle on $\mathbb{S}^{2}$ we refer to Figure 2.
@@
1: Draw $V\sim\mathcal{U}_{\mathbb{S}^{d-2}_{x}}$, call the result $v$.   {Perform Algorithm 5 with input xx.}
@@
3: Draw $\Theta\sim Q_{{x,}{v,}{t}}$, call the result $\theta$.   {Perform Algorithm 3 with input x,v,tx,v,t.}
@@
Observe that due to the $2\uppi$-periodicity of the great circle on the sphere, we have
@@
Proof (Proof of Theorem 16.) For arbitrary $x\in\mathbb{S}^{d-1}$, $s\in(0,\left\|p\right\|_{\infty})$, for any $v\in{\mathbb{S}^{d-2}_{x}}$ and $\theta\in[0,2\uppi)$ with $\gamma_{(x,v)}(\theta)\in L(s)$ we have (cf. (8)) that
@@
Moreover, for all $t\in(0,p(x))$ we obtain with $\Xi\sim\mathcal{U}_{(0,2\uppi)}$ that
$Q_{{x,}{v,}{t}}(B)\geq\mathbb{P}(\Xi\in B\cap L(x,v,t))=\frac{1}{2\uppi}{\lambda}(B\cap L(x,v,t)),\quad B\in\mathcal{B}([0,2\uppi)).$ (11)
@@
Note that the right hand-side of (9) coincides with the the right hand-side of (7) stated in Theorem 6 that addressed the uniform ergodicity statement for ideal slice sampling. This is due to the proof technique, since in the small set estimate, we lower bound both kernels with the same expression appearing in Lemma 11. Intuitively, it is clear that $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(H^{n}(x,\cdot),\pi)$ is smaller than $\sup_{x\in\mathbb{S}^{d-1}}d_{tv}(\widetilde{H}^{n}(x,\cdot),\pi)$, since the shrinkage procedure just adaptively imitates the acceptance/rejection step of the ideal one to gain computational efficiency. Exactly this gain in efficiency leads in applications to a potentially better accuracy to cost ratio, although the performance per Markov chain transition may be worse. We also point out that Remark 7 and Remark 8 apply to Theorem 16 as well.
@@ 4 Numerical Illustrations
We apply geodesic slice sampling on the sphere for approximate sampling on various distributions on the sphere and compare it to tailor-made and general purpose MCMC algorithms including random-walk Metropolis-Hastings (RWMH) and Hamiltonian Monte Carlo (HMC). The RWMH algorithm on the sphere is very similar to the standard RWMH, for details refer to Appendix F. Lan et al. (2014) have developed a modified version of HMC for sampling spherical distributions. The details of the HMC algorithm can be found in Appendix G. Both RWMH and HMC have a step-size parameter that is automatically tuned for each target, see Appendix H. For a prespecified sampler we denote by $(x_{n})_{n\in\mathbb{N}}\subset\mathbb{S}^{d-1}$ a realization of the corresponding Markov chain.
4.1 Rigid Registration of Biomolecular Structures
First, we discuss an application of the geodesic slice sampler on the sphere to a common inference problem in structural bioinformatics, computer vision, and robotics. Two 3D objects are represented by point clouds, $Q=\{\boldsymbol{q}_{i}\}_{i=1}^{I}$ and $P=\{\boldsymbol{p}_{j}\}_{j=1}^{J}$, where $\boldsymbol{q}_{i},\boldsymbol{p}_{j}\in\mathbb{R}^{3}$. The task is to find a 3D rotation matrix $\boldsymbol{R}\in SO(3)$ that best superimposes both point clouds. This is a rigid registration or pose estimation problem. In general, the point clouds have different size, and the correspondence between points in $Q$ and $P$ is unknown.
We use a probabilistic model inspired by the coherent point drift (CPD) method (Myronenko and Song, 2010) to address this task. According to the model, a point in $Q$ is either generated from a Gaussian mixture model whose components have equal weights and are centered on the rotated source $P$, or it is an outlier that has no corresponding point in $P$. Outliers are generated from a uniform distribution $\mathcal{U}_{B}$ over the bounding box $B$ of the target point cloud (i.e. $B\subset\mathbb{R}^{3}$ is the smallest rectangle such that $Q\subset B$). We denote the probability density function w.r.t. the Lebesgue measure of this distribution as $p_{o}$. The probability that a point is an outlier is $\omega\in[0,1]$. The probabilistic model, i.e., the conditional density again w.r.t. the Lebesgue measure, for a single point in $Q$ is
$p(\boldsymbol{q}_{i}\mid P,\boldsymbol{R},\omega,\sigma)=\omega p_{o}(\boldsymbol{q}_{i})+\frac{1-\omega}{J\,(2\pi\sigma^{2})^{3/2}}\sum_{j=1}^{J}\exp\left(-\frac{1}{2\sigma^{2}}\|\boldsymbol{q}_{i}-\boldsymbol{R}\boldsymbol{p}_{j}\|^{2}\right)\,.$ (12)
We can represent rotation matrices via unit quaternions (Horn, 1987) and thereby map the rigid registration problem to a distribution over $\mathbb{S}^{3}$. This mapping is achieved by representing rotation matrices with unit quaternions:
$\boldsymbol{R}(x)=\begin{pmatrix}1-2(x_{3}^{2}+x_{4}^{2})&2(x_{2}x_{3}-x_{1}x_{4})&2(x_{2}x_{4}+x_{1}x_{3})\\ 2(x_{2}x_{3}+x_{1}x_{4})&1-2(x_{2}^{2}+x_{4}^{2})&2(x_{3}x_{4}-x_{1}x_{2})\\ 2(x_{2}x_{4}-x_{1}x_{3})&2(x_{3}x_{4}+x_{1}x_{2})&1-2(x_{2}^{2}+x_{3}^{2})\end{pmatrix}$ (13)
where $x=(x_{1},x_{2},x_{3},x_{4})^{T}\in\mathbb{S}^{3}$. Therefore, probabilistic pose estimation boils down to generating samples from a distribution with density $p$ over $\mathbb{S}^{3}$, where $x\in\mathbb{S}^{3}$ encodes a 3D rotation matrix. The density $p$ w.r.t the volume measure on $\mathbb{S}^{3}$ of the overall posterior is (up to normalization) the product of the densities in (12), that is,
$p(x)\propto\prod_{i=1}^{I}p(\boldsymbol{q}_{i}\mid P,\boldsymbol{R}(x),\omega,\sigma)\,.$ (14)
In the following, we assume that the fraction of outliers $\omega$ and the standard deviation $\sigma$ are known such that the only unknown parameter is the unit quaternion $x$ encoding a rotation of the source relative to the target. We consider an example from structural bioinformatics and test MCMC samplers to tackle this rigid registration task. Adenylate kinase (AK) undergoes a large conformational change upon ligand binding. This internal structural change cannot be modeled as a global rigid transformation. Only the core domain, comprising a 60% subset of all points, can be related to each other via a rigid transformation. In our tests, the target and source point clouds show the closed and open conformation of AK (the protein data bank (PDB) codes 1AKE and 4AKE; see Fig. 3). We fix $\sigma=1$Å and $\omega=0.4$ (motivated by the fact that the core domain contains roughly 60% of all points); the bounding box volume of 1AKE is $5.73\times 10^{4}{\rm\AA }^{3}$. To guide the visual inspection, we use the chainbow coloring from blue to red offered by the Pymol software. Note that the posterior (14) is invariant under reordering of the points.
Figure 3: Rigid registration of point clouds. Each point cloud consists of 214 carbon-$\alpha$ atom positions. Points are colored according to their sequence position: N-terminal amino acids are shown in blue, C-terminal points are red. (A) Target point cloud showing AK in closed configuration. (B) Source point cloud showing AK in open configuration and in an orientation that differs from the orientation of the target. Note that target and source points with identical color correspond to each other. However, due to the non-rigid conformational change not all source points will generate target points. (C) Source rotated optimally so as to generate the target with high probability according to the CPD model (12). Note that the points in the source that do not generate a target point are shown in gray. These are located in the moveable domains of the AK structure (for example the orange domain that is closed in the target, but open in the source).
The posterior density (14) has many local peaks as is clear from Figure 4A showing a slice through $p$ along a random great circle passing through the posterior mode. Rigid registration by sampling from posterior (14) is challenging, because standard MCMC samplers tend to get stuck in subordinate peaks and miss the global maximum. To gain an overview of the posterior, we systematically evaluate $p$ over a regular tessellation of $\mathbb{S}^{3}$ based on the 600-cell (Straub et al., 2017). The 600-cell is the 4D analog of the regular icosahedron and can be used to approximate $\mathbb{S}^{3}$ at varying degrees of resolution. A single cell is a 3D tetrahedron whose corners lie in $\mathbb{S}^{3}$. By splitting each tetrahedron into eight equally-sized tetrahedra we obtain a tessellation at a higher resolution. Starting from the 600-cell encoding 300 unique rotations, we repeat the splitting process four times to obtain 1279264 unit quaternions that tessellate $\mathbb{S}^{3}$ in a regular fashion. Evaluation of $p$ at these quaternions offers an accurate discretization of the posterior that allows us to study its shape (see Figure 4B). The global maximum of $\log p$ on the quaternions of the 600-cell tessellation is 2192.89-2192.89. Although, an overwhelming fraction of quaternions achieves only very small posterior probabilities, almost all probability mass is concentrated in the maximum posterior peak: the fractional volume of $\mathbb{S}^{3}$ covered by quaternions $x\in\mathbb{S}^{3}$ with $\log p(x)>-2300$ is close to zero ($1.27\times 10^{-4}$ based on the 600-cell tessellation), but the cumulative probability is close to one ($1-4\times 10^{-44}$).
To investigate the robustness of probabilistic rigid registration with the geodesic slice samplers, RWMH and HMC, we launched 200 MCMC chains starting from random initial orientations that are generated by sampling unit quaternions uniformly from $\mathbb{S}^{3}$. We assess the performance of the samplers by counting how many among all 200 runs reached a log posterior probability greater than 2300-2300 and consider these simulations a “success”, because it located the dominant posterior mode rather than getting stuck in a subordinate peak. As shown in Figure 4C, the fraction of successful slice sampling runs increases with the number of MCMC iterations. As expected, the rejection strategy converges faster to the posterior mode than the shrinkage strategy, because it explores the sample space more exhaustively. The success rate of the other two MCMC samplers, RWMH and HMC, remains constantly low reaching only up to 7% after 2000 iterations. This demonstrates that geodesic slice sampling can deal with multi-modal posteriors, whereas RWMH and HMC tend to get stuck independent of the number of MCMC iterations. Both geodesic slice samplers reach a success rate of greater than 50% for about 50 iterations. The shrinkage strategy achieves 100% success rate after 1500 iterations, whereas the rejection strategy is at 100% success rate already after 200 iterations. In contrast, the performance of RWMH and HMC is quite poor with a success rates ranging between 3–7%. A similarly poor performance is also observed for a custom-made Gibbs sampler (not shown).
Rigid registration is often just one of many sampling tasks within the context of a more complex probabilistic model. Therefore, it is desirable that highly probable rotations are sampled after few MCMC iterations within a short computation time. This is not achieved by the rejection strategy (see Fig. 4D), because the dominant posterior mode is rather narrow resulting in many rejections. The shrinkage strategy, although requiring more steps to successfully detect the posterior peak with high probability, clearly outperforms any of the other samplers when taking also the computation time into consideration.
Figure 4: Probabilistic rigid registration of the open and closed structure of Adenylate kinase based on posterior $p$ given in (14). (A) Slice through $p$ along a great circle $x(\theta)=\cos\theta\,x_{1}+\sin\theta\,x_{2}$ where $x_{1}\in\mathbb{S}^{3}$ is the global optimum and $x_{2}\in{\mathbb{S}^{2}_{x_{1}}}$. (B) Marginal distribution of log posterior probabilities $\log p(x_{i})$ based on a discretization of $\mathbb{S}^{3}$ using $\sim 1.28\times 10^{6}$ regularly placed unit quaternions $x_{i}\in\mathbb{S}^{3}$. (C) Success rate against number of Markov chain iterations for all four samplers. (D) Success rate against computation time.
4.2 Mixture of von Mises-Fisher Distributions
To study the performance on a multi-modal target, we test the slice samplers as well as RWMH and HMC also on a $K$-component mixture model of von Mises-Fisher (vMF) distributions in $d$ dimensions. The vMF distribution is defined by the unnormalized density
$p_{\text{vMF}}{}(x;\mu,\kappa)=\exp(\kappa\mu^{T}\!x),\quad x\in\mathbb{S}^{d-1},$ (15)
where $\kappa>0$ is the concentration parameter and $\mu\in\mathbb{S}^{d-1}$. For a given $K\in\mathbb{N}$ our target distribution is a mixture of vMF distributions where each component has the same weight $1/K$ and a shared concentration parameter $\kappa$, i.e., the corresponding unnormalized density takes the form (note that the normalizers of each component are identical due to the shared concentration parameter)
$p_{\text{mix}}(x)=\sum_{k=1}^{K}p_{\text{vMF}}(x;\mu_{k},\kappa),\quad x\in\mathbb{S}^{d-1},$ (16)
where every $\mu_{k}\in\mathbb{S}^{d-1}$ with $k\in\{1,\dots,K\}$ is sampled w.r.t. the uniform distribution on the unit sphere and then fixed. We evaluate both geoSSS variants against RWMH and HMC on a $10$-dimensional mixture of vMF distributions with $K=5$ mixture components on the unit sphere $\mathbb{S}^{9}$. We set $\kappa=100$ and draw $10^{6}$ samples.
Figure 5: Marginal distributions (shown in colors) for the mixture of vMF distributions ($d=10,\kappa=100,K=5$) where columns corresponds to MCMC methods and rows correspond to the dimension of the marginal (only the first two dimensions are shown). Here, $\{e_{i}\}_{i=1}^{10}$ denotes the standard basis in $\mathbb{R}^{10}$. The black dashed line indicates the true marginals computed analytically for the mixture of vMF distribution.
To gain some insight into the performance of the samplers on the multi-modal target, we first look at the marginal distributions (see Fig. 5). We estimate marginal histograms from the samples and compare them with the marginal distributions of the vMF mixture that we computed analytically. Both variants of geoSSS represent the underlying mixture significantly better compared to HMC and especially RWMH which seems to explore only a single component of the vMF mixture.
Next we study quality measures to obtain more quantitative insights into the sampling performance. RWMH is the fastest to decorrelate, which is reflected by the rapid decay of the autocorrelation function (ACF) (see Fig. 6A). However, the reason for the fast decorrelation is that RWMH fails to explore the entire distribution, but is stuck in a single component. The geoSSS variants decorrelate much faster as compared to HMC. Note that the slow, linear decay of the ACF results from the fact that HMC spends a large fraction of the simulation time in a single mode and only occasionally escapes to another mode (also HMC only visits three out of all five modes in the course of the simulation).
Figure 6: (A) ACF plotted for the first dimension from the $10$-dimensional mixture of vMF distributions with $K=5$ and $\kappa=100$. (B) Kullback-Leibler divergence between the empirical and ideal frequency with which each mode is visited by the four MCMC samplers. (C) Geodesic distance (log-scale) between successive approximate samples.
As a measure for the quality of approximate sampling we consider the frequency with which the samplers visit the $K$ modes of the target mixture model (16). Ideally, these frequencies should be constant, because each component of the mixture has the same weight and concentration parameter. We contrast the empirical frequency $q_{k}$ with which the $k$-th mode is visited by a sampler with the uniform distribution by using the Kullback-Leibler (KL) divergence
$\text{KL}(q\mid\ell):=\sum_{k=1}^{K}q_{k}\log(q_{k}/\ell_{k})$
where $\ell=(\ell_{1},\ldots,\ell_{K})$ with $\ell_{k}=1/K$ for $k\in\{1,\ldots,K\}$ is the ideal distribution and $q=(q_{1},\ldots,q_{k})$. To estimate $q_{k}$, we assign each sample $x_{n}$ to the component under which $x_{n}$ is most likely, i.e.,
$k_{n}=\underset{k\in\{1,\ldots,K\}}{\text{argmax}}\,p_{\text{vMF}}(x_{n};\mu_{k},\kappa).$
Based on the allocations $k_{n}$, we can count how often each mode is visited by the Markov chain. A good MCMC sampler should produce $q_{k}\approx 1/K$ resulting in $\text{KL}(q\mid\ell)\approx 0$. The more the KL divergence differs from the minimum value of zero, the greater is the mismatch between the ideal and empirical frequency of mode visits. As seen from Fig. 6B, only the geoSSS variants achieve a small KL divergence. RWMH and HMC produce large KL divergences, indicating a complete failure to detect some of the components and, as a result, a poorer representation of the target distribution compared to geoSSS. Indeed, RWMH is stuck in a single mode in the course of the simulation. HMC fares better, but also misses two out of five modes.